OpenStudy (anonymous):

68 3/4 write as a fraction and round to the nearest hundredth

OpenStudy (anonymous):

\[\huge 68 \frac{3}{4} = \frac{68 \times 4 +3}{4}= \frac{272 +3}{4}\] \[\huge = \frac{275}{4}= 68.75\]

OpenStudy (anonymous):

\[\huge 68 \frac{3}{4} = 68 + \frac{3}{4} = 68+0.75 = 68.75\]

OpenStudy (timaashorty):

it's already written as a fraction @Nigil . Although I think you mean decimal.